# Muting Whisper: A Universal Acoustic Adversarial Attack on Speech Foundation Models

### v1

@@ 7 Conclusion
This work proposes a highly effective and practical method for ‘muting’ Whisper models, achieving a success rate of over 97%. A universal 0.64-second adversarial audio segment is trained to represent an acoustic realization of the <endoftext> token used by Whisper, such that when this audio segment is prepended to any speech signal, Whisper does not transcribe the speech, i.e., the model is ‘muted’. Further, this work finds that this universal acoustic adversarial segment transfer across different data distributions and can even transfer to different speech processing tasks. While this result offers a potential for speech privacy protection, it does also reveal the critical security implications of foundation models’ susceptibility to adversarial attacks. As speech processing systems continue to develop, addressing these vulnerabilities is an important direction for future research to ensure their reliability in real-world applications.
@@ 9 Risks and Ethics
10 Acknowledgements
This paper reports on research supported by EPSRC Project EP/V006223/1 (Multimodal Video Search by Examples) and Cambridge University Press & Assessment. We thank Mengie Qian for the preparation of the MGB dataset used in the experiments in this work.
@@ A.2 Models
Whisper model checkpoints are available in a range of sizes: Whisper tiny (39M parameters); Whisper base (74M); Whisper small (244M); Whisper medium (769M); and Whisper large (1.55B parameters). The Whisper models are available as English-only (en) or multilingual models. Whisper large is only available as a multilingual model. The Whisper models can be prompted to do speech recognition, voice activity detection, as well as speech translation and language identification for the multi-lingual model variants. This work considers a range of sizes of Whisper models for speech recognition and the multilingual versions are also evaluated for speech translation: tiny(.en), base(.en), small(.en) and medium(.en). The performance of each model, measured by the Word Error Rate (WER), for each dataset is given in Table 7.
@@ A.3 Attack Train Configuration
A further point to note is that we conducted separate experiments to confirm that when evaluating the adversarial attack, for no sample is the voice activity detector (used as part of Whisper’s transcription framework) returning ‘no speech’, i.e., the universal acoustic adversarial segment is a genuine realization of the <endoftext> special token. It is unlikely the voice activity detector would ever be activated at evaluation time as during the training of the universal attack segment the internal voice activity detector is not present.
@@ Appendix C Transferability Across Models
In this section we explore the transferability of the learnt universal acoustic adversarial attack segments across different Whisper models. Table 12 shows that there is no transferability across model sizes. This is an expected result: the attack aims to learn an acoustic realization of a non-acoustic event (<endoftext> token), but the fact that there is no real audio for the <endoftext> the attack is unable to find a genuine acoustic realization. Hence, the acoustic realization being learnt is specific realization of the <endoftext> token of the target model, such that the attack segment is specific to the parameter states of the target model. As different models have very different parameters and values, the learnt universal acoustic segments are distinct and do not transfer across the models. Nevertheless, the universal acoustic adversarial attack segment could always be trained with multiple target models and thus would be effective for multiple target models simultaneously. Future work will explore this notion further.

### v2

@@ 7 Conclusion
This work proposes a highly effective and practical method for ‘muting’ Whisper models, achieving a success rate of over 97%. A universal 0.64-second adversarial audio segment is trained to represent an acoustic realization of the <|endoftext|> token used by Whisper, such that when this audio segment is prepended to any speech signal, Whisper does not transcribe the speech, i.e., the model is ‘muted’. Moreover, this universal acoustic adversarial segment transfers across different data distributions and can even transfer to different speech processing tasks. While this result offers a potential for speech privacy protection, it does also reveal the critical security implications of foundation models’ susceptibility to adversarial attacks. As speech processing systems continue to develop, addressing these vulnerabilities is an important direction for future research.
@@ A.2 Models
Whisper model checkpoints are available in a range of sizes: Whisper tiny (39M parameters); Whisper base (74M); Whisper small (244M); Whisper medium (769M); and Whisper large (1.55B parameters). The Whisper models are available as English-only (en) or multilingual models. Whisper large is only available as a multilingual model. The Whisper models can be prompted to do speech recognition, voice activity detection, as well as speech translation and language identification for the multi-lingual model variants. This work considers a range of sizes of Whisper models for speech recognition and the multilingual versions are also evaluated for speech translation: tiny(.en), base(.en), small(.en) and medium(.en). The performance of each model, measured by the Word Error Rate (WER), for each dataset is given in Table 7. Further, in all experiments we use Whisper’s default decoding strategy with a beam size of 5.
@@ A.3 Attack Train Configuration
A further point to note is that we conducted separate experiments to confirm that when evaluating the adversarial attack, for no sample is the voice activity detector (used as part of Whisper’s transcription framework) returning ‘no speech’, i.e., the universal acoustic adversarial segment is a genuine realization of the <|endoftext|> special token. It is unlikely the voice activity detector would ever be activated at evaluation time as during the training of the universal attack segment the internal voice activity detector is not present.
@@ Appendix C Transferability Across Models
In this section we explore the transferability of the learnt universal acoustic adversarial attack segments across different Whisper models. Table 12 shows that there is no naive transferability of the adversarial audio segments across models. We next explain this result analytically. Based on the analysis, we further explore empirical methods to try and find adversarial audio segments that could transfer between models.
@@
C.1 Analytically understanding the transferability across models
Let $\mathbf{q}^{[k]}$ be the embedding generated by the final layer of the Transformer decoder, to be used to predict the next token (in the case of a muting whisper attack, the first token). For a vocabulary $\mathcal{V}$, we obtain the logits predicted by the model, $\mathbf{y}^{[|\mathcal{V}|]}$ via a projection matrix, $\mathbf{W}^{[|\mathcal{V}|\times k]}$  [^5] The projection matrix $\mathbf{W}$ is the same as the embedding matrix used at the input to the decoder.,
$\mathbf{y}=\mathbf{W}\mathbf{q},$ (10)
where a greedy decoder selects the token, $\hat{j}$ with the largest logit value,
$\hat{j}=\operatorname*{arg\,max}_{j}\{y_{j}\}.$ (11)
If we define the projection matrix using row vectors,
$\mathbf{W}=\begin{bmatrix}\text{-----}\mathbf{w}_{1}\text{-----}\\ \text{-----}\mathbf{w}_{2}\text{-----}\\ \vdots\\ \text{-----}\mathbf{w}_{|\mathcal{V}|}\text{-----}\end{bmatrix},$ (12)
then the greedily selected token can be equivalently selected as,
$\hat{j}=\operatorname*{arg\,max}_{j}\{\mathbf{w}_{j}^{T}\mathbf{q}\}$ (13)
If the first generated token is $\hat{j}=r$, then you would expect in a perfect system that the acoustic realization (audio segment), $\mathbf{x}_{r}$ of token $r$, when input to the encoder, to give,
$\mathbf{q}\approx\mathbf{w}_{r}$ (14)
to maximize its selection for generation. Note that you would expect that if row vectors $\mathbf{w}_{r}$ and $\mathbf{w}_{j}$ are geometrically close (cosine distance) (i.e. the predicted logit values $y_{r}$ and $y_{j}$ are positively correlated), the acoustic realizations $\mathbf{x}_{r}$ and $\mathbf{x}_{j}$ are similar too, i.e. token $r$ and token $j$ have a similar acoustic sound. We know that certain tokens have real acoustic sounds (that are model invariant), e.g., normal words like zoo, boy and hi have real acoustic realizations ($\mathbf{x}$) that are independent of specific models. Let $\mathcal{D}$ represent the set of tokens that have a real acoustic sound. Then for a model $\theta$, we can define the relative acoustic position of any token $r$ by considering its similarity to each of these real acoustic tokens.
$\bm{\rho}(r;\theta)=\left[\rho_{1}(r;\theta),\rho_{2}(r;\theta),\ldots,\rho_{|\mathcal{V}|}(r;\theta)\right]$ (15)
$\rho_{i}(r;\theta)=\begin{cases}\tilde{\mathbf{w}}_{i}^{T}\tilde{\mathbf{w}}_{r}&\text{if }i\in\mathcal{D}\\ \text{null}&\text{if }i\notin\mathcal{D}\end{cases}$ (16)
For a token $d\in\mathcal{D}$, where $\mathcal{D}$ represents all those tokens that have real sounds (e.g. normal words like hello, zoo, etc.), we would expect their relative positions (to other real sounds) to be very consistent across different models (This has been demonstrated in Table 13). If we define the difference in acoustic representation for any token $r$ as,
$s(r;\theta_{m},\theta_{n})=\|\bm{\rho}(r;\theta_{m})-\bm{\rho}(r;\theta_{n})\|_{2},$ (17)
then for a token $d\in\mathcal{D}$,
$\forall d\in\mathcal{D},\quad s(d;\theta_{m},\theta_{n})\leq\epsilon,$ (18)
where $\epsilon$ is an arbitrarily small value.
Token Model Top 5 closest tokens in $\mathcal{D}$ as per $\mathbf{W}$
zoo tiny Z, j, k, ch, iz
base Z, j, iz, ch, k
small Z, ch, iz, j, zh
medium Z, j, ch, rr, ez
boy tiny boys, girl, Boy, Bry, NOUN
base boys, girl, Boy, Missy, Cameraman
small boys, girl, Bry, Justin, Boy
medium boys, Boy, moil, ontec
hi tiny Hi, him, HI, iiii, high
base HI, Hi, iiii, Cameraman, Katie
small HI, Hi, pleasant, Julia, Hola
medium Hi, HI, FFFF, Adam, scream
eot tiny Male, Pro, Sa, Vict, Cho
base Arin, JIN, ELLE, ARRATOR, Jared
small WW, pleasant, Gra, Hyun, Missy
medium Everyone, sound, Something, Come, Aw
Table 13: Exploring the geometric relationship between embedding matrix tokens in $\mathbf{W}$. As expected, generally similar sounding words are close together. In some cases, similar domain/meaning words are also close. Note that there can be slight differences to the previous table as some tokens in the other table had a space before them.
The acoustic realization (sound) of the eot token is not known, such that $\texttt{eot}\notin\mathcal{D}$, as it’s not a real acoustic sound. However, we can predict which tokens the acoustic realization $\mathbf{x}_{\texttt{eot}}$ should be similar to, by considering the geometric position of $\mathbf{w}_{\texttt{eot}}$ relative to other tokens with a real sound, belonging to $\mathcal{D}$ we can compute $\bm{\rho}(\texttt{eot};\theta)$. For there to exist a muting adversarial attack audio segment, $\mathbf{x}_{\texttt{eot}}(\theta)$ that is transferable across different models, $\theta$, the acoustic realization of eot has to be the same/similar for the different models (the way the acoustic realization of any other real token in $\mathcal{D}$ is the same for all models). We can thus determine if there exists this true, universal acoustic realization of eot that is the same for all models by observing how consistent its relative position is to tokens in $\mathcal{D}$, as per $\bm{\rho}$. For a pair of models, $\theta_{m}$ and $\theta_{n}$, we expect there to exist a transferable muting adversarial attack, $\mathbf{x}_{\texttt{eot}}$ if,
$s(\texttt{eot};\theta_{m},\theta_{n})\leq\tau(\theta_{m},\theta_{n}),$ (19)
where we can define the threshold $\tau$ by considering the typical changes in similarity for other tokens with a real sound (belong to $\mathcal{D}$) that should have a consistent acoustic sound. We give error for variation by setting the threshold to be two standard deviations above the average change in similarity across models,
$\tau(\theta_{m},\theta_{n})=$ $\mathbb{E}_{r\in\mathcal{D}}[s(r;\theta_{m},\theta_{n})]$
$+2\cdot\sigma_{r\in\mathcal{D}}(s(r;\theta_{m},\theta_{n}))$ (20)
C.2 Empirical Evaluation of Model Transferability
We define the set of real acoustic sounds, $\mathcal{D}$ as the tokens which begin with any English letter (in roman alphabet) or English numeral (0-9). Table 14 reports uses the projection matrix, $\mathbf{W}$ of each Whisper model to determine the potential of the attack transferability. It is interesting to note that there is generally a low chance of model transferability, as the expected acoustic representation of the eot token is far less consistent than that of tokens with a real acoustic sound. These results demonstrate that there is no real audio representation for the <|endoftext|> token, and as a result the attack is unable to find a genuine acoustic realization. Hence, the acoustic realization being learnt is a specific realization of the <|endoftext|> token of the target model.
$\theta_{m}$ $\theta_{n}$ $s(\texttt{eot};\theta_{m},\theta_{n})$ $s(r;\theta_{m},\theta_{n})$
tiny.en base.en 12.13 $\underset{\pm 1.40}{3.50}$
tiny.en small.en 13.29 $\underset{\pm 1.80}{4.13}$
tiny.en medium.en 9.14 $\underset{\pm 1.37}{3.20}$
base.en small.en 19.72 $\underset{\pm 1.93}{5.61}$
base.en medium.en 6.65 $\underset{\pm 1.42}{4.32}$
small.en medium.en 13.40 $\underset{\pm 1.78}{3.81}$
tiny base 6.54 $\underset{\pm 0.35}{1.24}$
tiny small 4.57 $\underset{\pm 1.22}{5.46}$
tiny medium 4.21 $\underset{\pm 2.04}{6.16}$
base small 9.67 $\underset{\pm 1.19}{5.51}$
base medium 9.41 $\underset{\pm 2.12}{7.27}$
small medium 3.52 $\underset{\pm 1.57}{3.03}$
Table 14: Measuring theoretical potential transferability of muting attacks between models.
Nevertheless, next we explore methods to learn a universal attack segment that is able to transfer across the different models: we explicitly train the attack audio segment by considering multiple models at the same time during the training of the attack segment. We also explore initializing the attack audio segment with the optimal audio segments for single target models. The results are presented in Table 15. As expected from the above analysis, it is clear that it is difficult to learn an attack that can transfer across multiple models. However, we are able to obtain an audio attack segment that can transfer between the tiny and base model (when training to attack tiny, base and small), or between the tiny and medium models. Overall, this section has demonstrated that analytically there is little potential of a muting attack that can transfer between models because there is no real sound for the acoustic realization of the <|endoftext|> token, and therefore a specific acoustic realization is learnt for each specific target model.
Trn models Init eval model Performance
$\varnothing$ asl
tiny.en rand tiny.en 99.7 0.06
base.en 0.0 17.9
tiny.en, base.en rand tiny.en 99.42 0.160
base.en 0.00 17.79
base.en tiny.en 0.00 18.07
base.en 98.81 0.267
tiny.en, base.en, small.en rand tiny.en 98.43 0.590
base.en 99.12 0.227
small.en 0.00 17.75
small.en tiny.en 0.00 17.88
base.en 0.00 17.79
small.en 99.22 0.070
tiny.en, base.en, small.en, medium.en rand tiny.en 95.10 1.52
base.en 0.00 17.53
small.en 0.00 17.50
medium.en 98.33 0.670
Table 15: Training universal muting attack on multiple models. Training epochs is the maximum number of epochs required for each of the train (Trn) models when attacked individually. Initialization of the audio attack segment is either random or a previously targeted model. The best of 3 seeds is selected to obtain the most transferable attacks.